The diameter of 0---guage copper wire is 8.252 mm. Find the resistance of a 1.00-km length of such wire used for power transmission.

1
Expert's answer
2022-03-04T08:31:30-0500

Answer

Resistivity of copper wire

"\rho =1.77*10^{-8}\Omega-m"

Diameter

D

Length L=8.252mm

So resistence of this wire

"R=\rho \frac{A}{L}\\=\frac{1.77*10^{-8}*3.14D^2}{4*8.252*10^{-3}}\\=0.17*10^{-5}D^2 ohm"